OpenStudy (anonymous):

Chris went to the store to purchase 5.5 gallons of milk. When he got to the store, he noticed that the milk was no longer sold in gallons but in liters. Which of the following shows the number of liters, with the correct number of significant digits, Chris needs to buy in order to have 5.5 gallons of milk? (1 liter = 0.26418 gallon) A. 20.82 B. 21 C. 20.8 D. 20.819

OpenStudy (tanner23456):

make a proportion \[( \frac{ 1L }{ 0.26418 gallon })= (\frac{ 5.5L }{ x })\] rearrange the equaltion by cross multiplying x = 5.5 * 0.26418

OpenStudy (anonymous):

I got 1.45299

OpenStudy (tanner23456):

Sorry I know it's probably too late but I read the question wrong. I was in a hurry. It's 5.5 gallons not liters so the 5.5 should be on the bottom therefore it's 5.5/0.26418 = x and then you just choose the correct amount of Significant Figures.